A man of mass m moves with a constant speed on a plank of mass M ansd length L kept initially at rest on a frictionaless horizontal surface, from one end to the other in time t. the speed of plank relative to ground while man is moving is:
  1. L/t(M/m)
  2. L/t(m/(M+m))
  3. L/t(M/(M-m))
  4. none of these

Grade:11

2 Answers

Vikas TU
7 years ago
Let x be the distance covered by planck then,
Mx = m(L-x)
Solving for x,
(M+m)x = mL
x = mL / (M+m) .
Speed of planck = > x/t =  > mL / (M+m)t . 
i think none of these.
Annu
6 years ago
Let the distance moved by the plank be x.So the distance that the man will move is L - x.Applying Conservation of Linear Momentum,Mx = m(L - x)Mx = mL - mxMx + mx = mLx(M + m) = mLx = mL / (M + m)Distance moved by the plank in T seconds is mL/( M + m).Speed = Distance ÷ TimeSpeed = mL/ [ (M + m)T]